Encyclopedia Astronautica
2007.11.15 - ISS On-Orbit Status 11/15/07


Malenchenko and Whitson undertook the standard 30-min Shuttle RPM (R-bar Pitch Maneuver) skill training.

After wakeup and before breakfast, FE-2 Dan Tani again accessed the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and completing questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for later downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Peggy and Dan wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

FE-1 Yuri Malenchenko serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out" cycle to vacuum on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated at ~3:55pm EST. (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. Filter bed 1 was regenerated yesterday.)

Malenchenko & Whitson undertook the standard 30-min Shuttle RPM (R-bar Pitch Maneuver) skill training, Peggy's first, Yuri's second, using DCS-760 digital still cameras with 400 & 800mm lenses at Service Module (SM) windows 6 & 8 to take imagery of documented CEO Earth targets facing the velocity vector (i.e., in flight direction). Afterwards, Peggy downlinked the obtained images to the ground for analysis, to be discussed at a subsequent tagup. (The skill training prepares crewmembers for the bottomside mapping of the Orbiter at the arrival of STS-122/1E in December. During the RPM at ~600 ft from the station, the ISS crew will have only ~90 seconds for taking high-resolution digital photographs of all tile areas and door seals on the Atlantis from SM windows 6 & 8, to be downlinked for launch debris assessment. Thus, time available for the shooting will be very limited, requiring great coordination between the two headset-equipped photographers and the Shuttle.)

The FE-2 worked on the EMCS (European Modular Cultivation System) facility, removing MULTIGEN-1 (Molecular & Plant Physiological Analyses of the Microgravity Effects on Multigenerations of Arabidopsis thaliana) experiment containers from the rotors, as MULTIGEN-1 has reached the end of its 75-day run. (MULTIGEN was activated on 8/25. This ESA experiment evaluates the morphology of A. thaliana over three generations. The plants grown in EMCS are the First Generation of A. thaliana to be grown on ISS. A. thaliana, commonly know as Thale Cress, is a model organism to study genetic morphological events since its genome has been mapped and is widely used in genetic and molecular research. MULTIGEN studies the morphological and genetic effects of microgravity on plants, which will be grown to maturity (7-8 weeks) and then dried and stored for return to Earth. Some of the seeds that are returned will be used in ground studies and others will be returned to ISS for the Second Generation of growth. The Second Generation's seeds will be analyzed genetically and also used for a Third Generation of plants on ISS. Ultimately, the data collected from this multigenerational experiment could be applied to future agriculture endeavors for long-duration space flight.)

The CDR and FE-2 spent ~40 min in the U.S. Airlock (A/L), successfully completing the standard water recharge of the EMUs (Extravehicular Mobility Units) used by her and Yuri during the recent EVA-10 'Alpha'. (The MIB (Mishap Investigation Board) that evaluated the results from the chamber EMU inspection -- after a 'smell of smoke' was noticed in a SSATA (Space Station Airlock Test Article) chamber on 11/9 -- has cleared the on-orbit suits for the normal EMU water recharge (the original Flight Rule limit of 48 hrs for post-EVA recharge was extended to 7 days). The EMUs have the Go for spacewalk use.)

Afterwards, the crew had two hours for reviewing the latest uplinked EVA procedures material and crib sheets, particularly on ammonia decontamination, of the upcoming spacewalk, EVA-11 'Bravo' on 11/20 (next Tuesday), followed by a teleconference with ground personnel to discuss details.

(After the standard A/L Campout Monday night, EVA-11 is expected to last about 6h 40m. Its Node-2 Outfitting objectives are:

Remove Stbd NH3 (ammonia) Shunt Jumper;
Configure Vent Tools;
Vent & Stow the Stbd NH3 Shunt Jumper;
Release Node-2 Fluid Caps, reconfigure P1 radiator SFUs (Squib Firing Units, fired today), and release Node-2 Loop A Fluid Tray;
Relocate Node-2 Loop A Fluid Tray, then attach it, deploy/mate it, and vent N2 (nitrogen) from it;
Mate & open hinge QDs (Quick Disconnects), S0 Fluid QDs, and Node-2 Fluid QDs;
Connect Node-2 Fluid Line Heater Cables;
Connect Node-2 Port Avionics to Node-2; and
Mate Primary PMA-2/Node-2 Umbilicals.

Potential get-ahead tasks are:

Relocate N2 VTE (Vent Tool Extender) bag for EVA-12 'Charlie';
Relocate APFR (Articulating Portable Foot Restraint) for EVA-12;
Connect Stbd Avionics umbilicals to Node-2;
Connect PMA-2 redundant umbilicals;
Deploy SSPTS (Station-to-Shuttle Power Transfer System) cable; and
Relocate APFR for 1E.

Continuing the interior outfitting of Node-2, Peggy Whitson completed a major task conducting activation & checkout of SPS ELPS (Secondary Power System/Emergency Light Power Supply) systems in Harmony (22 bolts per panel!).

The CDR also performed the periodic thorough inspection of ELPS units in Node-1. (There are three ELPS units in the Node.)

The FE-2 meanwhile inspected and checked out the ELPS in the A/L, which has one unit.

In Node-2, Dan Tani removed the PPRV (Positive Pressure Relief Valve), now no longer required, and installed an MPEV (Manual Pressure Equalization Valve) in its place.

Whitson conducted the weekly 10-min. CWC (Contingency Water Container) audit as part of the continuing WDS (Water Delivery System) assessment of onboard water supplies. (Updated 'cue cards' based on the crew's water calldowns are sent up every other week. Preliminary analysis of samples of water transferred in CWCs from Discovery to ISS during 10A indicates bacterial contamination in all containers transferred. This newly transferred water (no more critical than common faucet water on the ground) is not scheduled to be used in the near future, and there are ample other supplies on board. An investigation is underway to develop a forward plan.)


FE-1 Malenchenko set up the pumping equipment and initiated (later closed out) the periodic transfer of urine from 4 EDV-U containers in the SM to the Rodnik BV2 tanks of Progress M-61 (26P). (Leak checks performed by the crew on 10/8-9 on the membrane (expulsion bladder) of the Progress' Rodnik BV1 tank showed that BV1 is unsuitable for liquid waste transfer due to a small leak in the tank's bladder.)

Dan Tani took air samples for the periodic (weekly) atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the hand-held CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit). Batteries are to be replaced if necessary. (Purpose of the 10-min. activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements.)

Peggy Whitson connected the regular ITCS LTL (Internal Thermal Control System/Low Temperature Loop) coolant jumper to the LAB1D6 rack, to support the ground-commanded activation of the U.S. CDRA (Carbon Dioxide Removal Assembly), and Houston lowered the temperature setpoint to the regular 9.4 degC. (CDRA activation is from 1:31pm-2:31pm.)

Afterwards, the CDR activated the AR SDS (Atmosphere Revitalization/Sample Delivery System) in Node-2 by opening the SDS manual valves to allow for MCA (Major Constituents Analyzer) sampling of Node-2. (The SDS, along with the MCA, PCA (Pressure Control Assembly), TCCS (Trace Contaminant Control Subassembly) and CVV (Carbon Dioxide Vent Valve assembly), is a subsystem of the Atmosphere Control & Supply System of the Lab's ECLSS (Environment Control & Life Support System).)

Yuri completed the routine maintenance of the SOZh system (Russian ECLSS) in the SM, including ASU toilet facilities systems/replaceables; and

Dan filled out the regular FFQ (Food Frequency Questionnaire), his second, on the MEC (Medical Equipment Computer). (By means of these FFQs, U.S. astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins.)

The FE-2 also conducted the periodic (every two weeks) inspection of the RED (Resistive Exercise Device) canister cords and accessory straps as well as the canister bolts for re-tightening if required.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Tani transferred the crew's exercise data file to the MEC for downlink.

Dan had another 60 minutes for himself for general orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

At ~3:15am EST this morning, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.

At ~4:05am, Yuri linked up with TsUP stowage specialists via S-band to conduct the weekly IMS (Inventory Management System) tagup, discussing stowage issues and equipment locations. (Issues of discussion today pertained to updating equipment locations assigned to Soyuz TMA-11, including packing of return equipment already stowed on ISS.)

Two items waiting for Whitson and Tani on their discretionary 'job jar' task list, besides the SLEEP logging (see above), are two audits - of onboard CTBs (Cargo Transfer Bags) and the 'stowage rack of the week', currently the one at LAB1P5 -- using a new Audit function introduced with the implementation of the laptop-based IMS 2.0 version on Increment 15. (This function allows the user to setup audits of bags, kits, stowage locations, just about anything that is considered a container, via a special Audit window with a linked Object Properties window.)

P1 Radiator Deployment: This morning at ~6:00-7:00am EST, the ground successfully deployed the ETCS (External Thermal Control System) Loop B PVRs (Photovoltaic Radiators) #1 & #3 on the P1 truss, observed by video camera on the SSRMS robotarm, after it ungrappled Node-2 but remains based on the Lab PDGF (Power & Data Grapple Fixture). (Russian thrusters were inhibited from 5:10-7:32am, and exercise prohibited during the deployment. Following the deployment, ETCS Loop A & B temperature setpoints were raised from 2.8 to 6.1 degC in preparation for commissioning of the six Node-2 IFHX's (Interface Heat Exchangers) during EVAs 11 (Bravo) and 12 (Charlie).)

CEO (Crew Earth Observation) photo targets uplinked for today were Ouarkziz Impact Crater (Ouarkziz is a relatively "young" impact crater, dated at 70 million years. The crater is 3.5 km in diameter and is located in the Atlas Mountains of western Algeria near the border with Morocco. As ISS crossed the Atlas Mountains from the NW in late morning, the crew was to shoot left of track for the impact in folded rocks near the south side of the E-W mountain range), S. Mozambique (this target area is in the coastal plains of far southern Mozambique just north of the capital city of Maputo. Rapid development of this area is now being planned. To better monitor this change, CEO observers are seeking baseline, pre-development imagery. ISS approach was from the NW just after midday. Fair weather was expected and observers requested a nadir mapping strip of this area along the orbit track), and Patagonian Glaciers (the Northern Patagonian Ice Field is the northernmost of two large remaining ice fields of South America, located in the Andes Mountains of southern Chile. It is also the world's largest ice field so far from the poles. The crew had a near-nadir, fair-weather pass in mid-afternoon. They were to try for detailed views of the glacier features, particularly on the western and northern margins.)

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use